OpenStudy (anonymous):

extremely confused. Algebra 2 Word problem. At a local produce stand sugar-coated strawberries sell for $3 dollare per pund and chocolate covered bananas sell for $4 per ound. How many punds of strawberries must be purchased to have a 5-pound mixture that costs $3.60 per pound?

OpenStudy (anonymous):

Set up a system of two linear equations: one for total weight, one for total cost.

OpenStudy (anonymous):

e.g. You can use S=pounds of strawberries, and B=pounds of bananas, so S+B=5

OpenStudy (anonymous):

Can you get the second equation for total cost?

OpenStudy (anonymous):

4b + 3s = 3.60 ?

OpenStudy (anonymous):

Very close! The $3.60 is per pound. You need to multiply that by 5 pounds to get the total cost.

OpenStudy (anonymous):

so 4b+3s=18 ?

OpenStudy (anonymous):

That should do it.

OpenStudy (anonymous):

yah im stuck I got to s = 5,75 - 5b/4

OpenStudy (anonymous):

this looks like a good idea, but may be too much work if you call the number of pounds of bananas say \(x\) then since the number of pounds total is 5 the number of pounds of strawberries is necessarily \(5-x\) since they have to add to 5 then the total cost will be \[4x+3(5-x)\] and you want that to be \(3.6\times 5=18\) so solve \[4x+3(5-x)=18\] for \(x\)

OpenStudy (anonymous):

in other words take your first equations S+B=5 solve for B

OpenStudy (anonymous):

sub into the second equation

OpenStudy (anonymous):

@satellite73 i got x= 3 now what

OpenStudy (anonymous):

If the total weight is 5 pounds and you have 3 pounds of bananas . . .

OpenStudy (anonymous):

so 2 pounds of strawberries? right

OpenStudy (anonymous):

You can check it by putting those numbers back into your original equations to see if they both work.

OpenStudy (anonymous):

that is why it is very very important to write as a first step what your variable represents notice the first think i wrote was "call the number of pounds of bananas say \(x"\)

OpenStudy (anonymous):

that is so when you solve, you can go back and say "oh, what was \(x\) i forget?" and then look and see that \(x\) is a number, namely the number of pounds of bananas in this case so you can be assured that you have 3 pounds of bananas (in this example)

OpenStudy (anonymous):

if you are still there i hope i have convinced you that you should not skip this step, actually write down on paper "let \(x\) be the number of ..."

OpenStudy (anonymous):

@satellite, that's why I recommended S=pounds of strawberries, and B=pounds of bananas; not only defines the variables, but they are easy to remember which is which.

OpenStudy (anonymous):

yeah that is actually a better plan, isn't it? i am just so used to typing \(x\) that when i try with other variables i make lots of typos.

OpenStudy (anonymous):

I know what you mean. Sometimes, good ol' x and y make everything familiar and easy to work with. ;-)